# Enhancing Wrist Abnormality Detection with YOLO: Analysis of State-of-the-art Single-stage Detection Models

### v1

@@ Abstract
Diagnosing and treating abnormalities in the wrist, specifically distal radius, and ulna fractures, is a crucial concern among children, adolescents, and young adults, with a higher incidence rate during puberty. However, the scarcity of radiologists and the lack of specialized training among medical professionals pose a significant risk to patient care. This problem is further exacerbated by the rising number of imaging studies and limited access to specialist reporting in certain regions. This highlights the need for innovative solutions to improve the diagnosis and treatment of wrist abnormalities. Automated wrist fracture detection using object detection has shown potential, but current studies mainly use two-stage detection methods with limited evidence for single-stage effectiveness. This study employs state-of-the-art single-stage deep neural network-based detection models YOLOv5, YOLOv6, YOLOv7, and YOLOv8 to detect wrist abnormalities. Through extensive experimentation, we found that these YOLO models outperform the commonly used two-stage detection algorithm, Faster R-CNN, in bone fracture detection. Additionally, compound-scaled variants of each YOLO model were compared, with YOLOv8x demonstrating a fracture detection mean average precision (mAP) of 0.95 and an overall mAP of 0.77 on the GRAZPEDWRI-DX pediatric wrist dataset, highlighting the potential of single-stage models for enhancing pediatric wrist imaging.
@@ 1 Introduction
Wrist abnormalities are a common occurrence in children, adolescents, and young adults. Among them, wrist fractures such as distal radius and ulna fractures are the most common with incidence peaks during puberty Hedstrom et al. (2010); Randsborg et al. (2013); Landin (1997); Cheng and Shen (1993). Timely evaluation and treatment of these fractures are essential to prevent life-long implications. Digital radiography is a widely used imaging modality to obtain wrist radiographs which are then interpreted by surgeons or physicians in training to diagnose wrist abnormalities. However, medical professionals may lack the specialized training to assess these injuries accurately and may rely on radiograph interpretation without the support of an expert radiologist or qualified colleagues Hallas and Ellingsen (2006). Studies have shown that diagnostic errors in reading emergency X-rays can reach up to 26% Guly (2001); Mounts et al. (2011); Er et al. (2013); Juhl et al. (1990). This is compounded by the shortage of radiologists even in developed countries Burki (2018); Rimmer (2017); Makary and Takacs (2022) and limited access to specialist reporting in other parts of the world Rosman (2015) posing a high risk to patient care. The shortage is expected to escalate in the upcoming years due to a growing disparity between the increasing demand for imaging studies and the limited supply of radiology residency positions. The number of imaging studies rises by an average of five percent annually, while the number of radiology residency positions only grows by two percent. Smith-Bindman et al. (2019). While imaging modalities such as MRI, CT, and ultrasound can assist in the diagnosis of wrist abnormalities, some fractures may still be occult Fotiadou et al. (2011); Welling et al. (2008); Neubauer et al. (2016).
Recent advances in computer vision, more specifically, object detection have shown promising results in medical settings. Some of the positive results of detecting pathologies in trauma X-rays were recently published Adams et al. (2020); Tanzi et al. (2020); Chung et al. (2018); Choi et al. (2020). Computer vision algorithms are accurate, efficient, and more importantly extremely quick to produce results compared to any radiologist or other imaging modalities currently used in practice. For example, radiology imaging delays have been found to independently contribute to longer hospital stays, as indicated by a recent study Courane et al. (2016). In addition, a separate study Perotte et al. (2018) found that creating reports from CT scans often took over three hours, with radiologists being responsible for a significant portion (42%) of the delay. The delays in obtaining clinically relevant information can have significant impacts on patients and contribute to unnecessary burdens on health systems, patients, and insurers. Computer vision algorithms can potentially address the delays associated with radiographic interpretation by providing a more efficient and prompt alternative, while still achieving comparable or even superior results.
@@
Object detection has emerged as a powerful tool for identifying abnormalities in X-ray images. Its ability to locate and classify various objects within an image has made it a valuable asset in the diagnosis and treatment of various medical conditions. In recent years, significant progress has been made in the development of object detection algorithms, leading to their widespread adoption in the medical community. An earlier approach called the sliding window approach Lampert et al. (2008) for object detection involved dividing an image into a grid of overlapping regions and then classifying each region as containing the object of interest or not. There are several disadvantages of this approach, one of them being that it is computationally expensive as a large number of regions need to be classified. To address these issues, region-based methods were invented. The introduction of Region-based Convolutional Neural Network (R-CNN) Girshick et al. (2013b) was the first breakthrough in the application of region-based methods. The main idea behind these methods was to generate candidate object regions and classify only those regions as containing the object of interest or not.
@@
Another method developed as an improvement over the sliding window approach was the single-stage detection method which has gained popularity in recent years due to its efficiency and good performance. This approach uses a single forward propagation through the network to predict bounding boxes and class probabilities, eliminating the need to generate candidate object regions, and making it faster than region-based approaches. While two-stage detection generates candidate regions in the first stage and refines them in the second stage at the cost of speed, single-stage detection provides a balance between speed and accuracy by predicting final results in a single pass through the network.
@@
Two-stage detection has been the most widely used approach for detecting wrist abnormalities in recent years. However, there has been limited research on the effectiveness of single-stage detectors in detecting various abnormalities in the wrist, including fractures. In this study, we focus on the effectiveness of single-stage detectors in detecting wrist abnormalities, more specifically, we focus on the capabilities of recent versions of the YOLO algorithm (v5, v6, v7, and v8). Additionally, this study is unique in its use of a large, comprehensively annotated dataset called GRAZPEDWRI-DX presented in a recent publication Nagy et al. (2022). The characteristics and complexity of the dataset are discussed in section 4
@@ 1.1 Study Objective & Research Questions
The primary objective of this study is to test the effectiveness of the state-of-the-art YOLO detection models, YOLOv5, YOLOv6, YOLOv7, and YOLOv8 on a comprehensively annotated dataset "GRAZPEDWRI-DX" Nagy et al. (2022) recently released to the public. We compare the performances of all variants within each YOLO model employed in this study to see whether the use of a compound-scaled version of the same architecture improves its performance. Moreover, this study also investigates how effective these single-stage detection methods are in detecting fractures compared to a two-stage detection method widely used in the past. We hypothesize that fractures in the near vicinity of the wrist in pediatric X-ray images can be detected efficiently using YOLOv5, YOLOv6, YOLOv7, and YOLOv8 models proposed by ultralytics (2022), Li et al. (2022), Wang et al. (2022), and ultralytics (2023) respectively. We prove our hypothesis using the comprehensively annotated GRAZPEDWRI-DX dataset.
The general objective of the study is to use the GRAZPEDWRI-DX dataset to analyze the potential of utilizing object detection techniques in answering the following research questions (RQ):
@@
    In the analysis of wrist images, do the state-of-the-art single-stage detection models outperform a two-stage detection model widely used in the past?
@@
    Does the use of compound scaled variants within each YOLO algorithm improve its performance in detecting fractures in the vicinity of the wrist in pediatric X-ray images?
@@ 1.2 Contribution
  • A thorough performance assessment of state-of-the-art YOLO detection models (YOLOv5, YOLOv6, YOLOv7, and YOLOv8) on the newly released GRAZ-PEDWRI-DX dataset, a large and diverse set of pediatric X-ray images. To the best of our knowledge, this is the first study of its kind.
  • An in-depth comparison of the performance of various variants within each YOLO model utilized, including compound, medium, and smaller-scaled versions.
  • Achieved state-of-the-art mean average precision (mAP) score on the (GRAZPEDWRI-DX dataset).
  • A detailed analysis of the performance of single-stage detection models in comparison to the two-stage detection model widely used in the literature, namely, Faster R-CNN.
@@ 3.1 Research Design
3.2 Study Dimensions
The following dimensions are used to facilitate the interpretation of results:
  • Abnormality-(ab): The object detection models were evaluated on their ability to detect different types of abnormalities in the radiographic images.
  • Fracture-(f): The object detection models were also evaluated on their ability to effectively detect fractures in the radiographic images.
  • Recall-(r): The proportion of positive instances that were correctly detected by the model. The calculation of recall is TP / (TP + FN), where TP represents the number of true positive detections and FN the number of false negative detections.
  • Precision-(p): The proportion of positive detections that were actually positive instances. It is calculated by dividing the number of true positive detections (TP) by the sum of true positives and false positives (incorrect detections) represented as TP / (TP + FP), where FP stands for false positive detections.
  • Mean Average Precision-(mAP): mAP is a performance metric used to evaluate an object detection model with an intersection over union (IoU) threshold of 0.5. It’s a widely adopted evaluation method for object detection models as it takes into account both precision and recall.
@@ 3.4 Deep Learning Models For Object Detection
In this study, we employed 4 single-stage detection models, namely YOLOv5, YOLOv6, YOLOv7, and YOLOv8, as well as a two-stage detection model Faster R-CNN. To further optimize the performance of the single-stage models, we experimented with multiple variations of each YOLO model, ranging from 5 to 7 variants. This resulted in a total of 23 wrist abnormality detection procedures.
We conducted initial training on various variants of each YOLO model. Subsequently, we selected the highest-performing variant within each YOLO model based on the results obtained and compared them to the two-stage detection model Faster R-CNN. The models that were trained for 100 epochs, were observed to converge between 90-100 epochs, indicating no additional improvement beyond the 100th epoch, thus further training was deemed unnecessary.
The YOLO (You Only Look Once) algorithm, initially introduced by Redmon et al. (2015) in 2016, is a single-stage object detection approach that uses a single pass of a convolutional neural network (CNN) to make predictions about the locations of objects in an image, making it faster than other approaches to date. In 2021, YOLOv4 achieved the highest mean average precision on the MS COCO dataset while also being the fastest real-time object detection algorithm Bochkovskiy et al. (2020). Since its initial release, the algorithm has undergone several improvements, with versions ranging from v1 to v7 Redmon et al. (2015); Redmon and Farhadi (2016); Farhadi and Redmon (2018); Bochkovskiy et al. (2020); Jiang et al. (2022), with v5, v6, v7, and v8 being released in 2020, 2022, and 2023 offering smaller volume, higher speed, and higher precision Solawetz (2020); Nelson (2020); Jiang et al. (2022). Fig. 2 illustrates the general structure of YOLO with backbones used in this study such as CSP, VGG, and EELAN.
@@
R-CNN proposed by Girshick et al. (2013a) was one of the first algorithms to achieve state-of-the-art performance on the PASCAL VOC object detection benchmark. R-CNN is a two-stage algorithm that takes an entire image as input, generates regions likely containing objects, extracts features using a CNN, and classifies objects within these regions. Faster R-CNN is a widely adopted and well-established model within the R-CNN family, known for its efficiency and accuracy in object detection. It has been widely utilized in the medical field, specifically in the detection of bone fractures. The Faster R-CNN model, first introduced by Ren et al. (2015), has continued to be a significant and influential contribution to the field of computer vision, remaining one of the most highly cited papers in the field to this day.
@@ 3.4.1 The YOLOv5 Model
The YOLO framework consists of three main components: the backbone, the neck, and the head. First, the input terminal performs various data processing tasks, including adaptive image filling and mosaic data augmentation Solawetz (2020). In our research, we have utilized the same data augmentation and pre-processing methods. Additionally, the YOLOv5 model uses adaptive anchor frame calculation to optimize its performance on different datasets by adjusting its anchor frame size when the dataset changes.
The backbone is responsible for extracting image features. It aggregates and forms features at different granularities. The specific CNN architecture used in YOLOv5 is CSPDarknet since this is the best-performing one so far Wang et al. (2020). Hence, in our work, we have utilized the same CNN architecture. The CSPDarknet architecture consists of convolutional, pooling, and residual connections represented mathematically as:
@@
where $F_{i}$ is the feature maps at the $i$-th layer, $F_{i-1}$ is the feature maps at the $(i-1)$-th layer, $W_{i}$ are the weights and biases at the $i$-th layer, and $f(\cdot)$ is the function applying convolution and pooling operations. The SPP structure is then applied to the feature maps produced by the CSPDarknet to extract features at multiple scales. This can be represented mathematically as:
@@
where $F_{SPP}$ represents the multi-scale feature maps produced by the SPP structure, and $g(\cdot)$ represents the function that applies the SPP operation to the input feature maps $F_{i}$.
The neck of YOLOv5 uses Path Aggregation Network (PANet) to aggregate features from the backbone and produce higher-level features for the output layers. The same architecture is used in our study. The head constructs output vectors with class probabilities, objectness scores, and the bounding box (coordinates for the box: center, height, and width) representing objects in the image. The utilized head is the same as the original implementation.
The YOLOv5 model includes five different model variants, namely, YOLOv5n, YOLOv5s, YOLOv5m, YOLOv5l, and YOLOv5x. All are compound-scaled variants of the same architecture. Each of these offers different detection accuracy and performance. The differences between these variants are the number of channels i.e. (the depth of the network), and the number of layers.
Figure 3: Performance comparison of YOLOv5 variants on COCO ultralytics (2022).
In the experimentation of YOLOv5 variants, standard hyperparameters were utilized. The input resolution was fixed at 640 pixels, and the batch size was set to 16. The optimization algorithm employed was Stochastic Gradient Descent (SGD) with an initial learning rate $\alpha=1\times 10^{-2}$, final learning rate $\alpha_{f}=1\times 10^{-2}$, momentum = 0.937, weight decay = $5\times 10^{-4}$, warmup momentum = 0.8, and warmup bias lr = 0.1. Each variant underwent 100 epochs of training from scratch. During the evaluation phase, each variant was tested on 1016 randomly selected samples, using an Intersection over Union (IoU) threshold of 0.5 for inference.
@@ 3.4.2 The YOLOv6 Model
YOLOv6 features an anchor-free design and reparameterized Backbone, with VGG and CSP Backbones used in the "n" and "s" variants, and "m", "l" and "l6" variants respectively. This Backbone is referred to as EfficientRep. The Neck, named Rep-PAN, is similar to YOLOv5, but the Head is efficiently decoupled, improving accuracy and reducing computation by not sharing parameters between the classification and detection branches. The YOLOv6 also includes five different model variants, namely, YOLOv6n, YOLOv6s, YOLOv6m, YOLOv6l, and YOLOv6l6. Fig. 4 shows the performance comparison of these variants on the COCO dataset.
Figure 4: Performance comparison of YOLOv6 variants on COCO meituan (2023).
All 5 variants of YOLOv6 were trained for 100 epochs from scratch. Standard hyperparameters were used. The input was set to 640 pixels with a batch size of 16 samples. The optimization algorithm employed was Stochastic Gradient Descent (SGD) with the same parameters as used for YOLOv5, including initial and final learning rates, momentum, warmup momentum, weight decay, and warmup bias lr. As before, each variant was tested on 1016 randomly selected samples, using an Intersection over Union (IoU) threshold of 0.5 for inference.

### v2

@@ Abstract
Diagnosing and treating abnormalities in the wrist, specifically distal radius, and ulna fractures, is a crucial concern among children, adolescents, and young adults, with a higher incidence rate during puberty. However, the scarcity of radiologists and the lack of specialized training among medical professionals pose a significant risk to patient care. This problem is further exacerbated by the rising number of imaging studies and limited access to specialist reporting in certain regions. This highlights the need for innovative solutions to improve the diagnosis and treatment of wrist abnormalities. Automated wrist fracture detection using object detection has shown potential, but current studies mainly use two-stage detection methods with limited evidence for single-stage effectiveness. This study employs state-of-the-art single-stage deep neural network-based detection models YOLOv5, YOLOv6, YOLOv7, and YOLOv8 to detect wrist abnormalities. Through extensive experimentation, we found that these YOLO models outperform the commonly used two-stage detection algorithm, Faster R-CNN, in fracture detection. Additionally, compound-scaled variants of each YOLO model were compared, with YOLOv8m demonstrating a highest fracture detection sensitivity of 0.92 and mean average precision (mAP) of 0.95. On the other hand, YOLOv6m achieved the highest sensitivity across all classes at 0.83. Meanwhile, YOLOv8x recorded the highest mAP of 0.77 for all classes on the GRAZPEDWRI-DX pediatric wrist dataset, highlighting the potential of single-stage models for enhancing pediatric wrist imaging.
@@ 1 Introduction
Wrist abnormalities are a common occurrence in children, adolescents, and young adults. Among them, wrist fractures such as distal radius and ulna fractures are the most common with incidence peaks during puberty Hedstrom et al. (2010); Randsborg et al. (2013); Landin (1997); Cheng and Shen (1993). Timely evaluation and treatment of these fractures are essential to prevent life-long implications. Digital radiography is a widely used imaging modality to obtain wrist radiographs. While X-ray is often the first and most common imaging modality used for wrist problems, the choice of test depends on the suspected abnormality, clinical presentation, and available resources. If an X-ray doesn’t provide a clear diagnosis, other imaging modalities like MRI, CT, or ultrasound may be recommended. The obtained radiographs are then interpreted by surgeons or physicians in training to diagnose wrist abnormalities. However, medical professionals may lack the specialized training to assess these injuries accurately and may rely on radiograph interpretation without the support of an expert radiologist or qualified colleagues Hallas and Ellingsen (2006). Studies have shown that diagnostic errors in reading emergency X-rays can reach up to 26% Guly (2001); Mounts et al. (2011); Er et al. (2013); Juhl et al. (1990). This is compounded by the shortage of radiologists even in developed countries Burki (2018); Rimmer (2017); Makary and Takacs (2022) and limited access to specialist reporting in other parts of the world Rosman (2015) posing a high risk to patient care. The shortage is expected to escalate in the upcoming years due to a growing disparity between the increasing demand for imaging studies and the limited supply of radiology residency positions. The number of imaging studies rises by an average of five percent annually, while the number of radiology residency positions only grows by two percent. Smith-Bindman et al. (2019). While imaging modalities such as MRI, CT, and ultrasound can further assist in the diagnosis of wrist abnormalities, some fractures may still be occult Fotiadou et al. (2011); Neubauer et al. (2016).
@@
Recent advances in computer vision, more specifically, object detection have shown promising results in medical settings. Some of the positive results of detecting pathologies in trauma X-rays were recently published Adams et al. (2020); Tanzi et al. (2020); Choi et al. (2020). In recent years, significant progress has been made in the development of object detection algorithms, leading to their widespread adoption in the medical community. An earlier approach called the sliding window approach Lampert et al. (2008) for object detection involved dividing an image into a grid of overlapping regions and then classifying each region as containing the object of interest or not. Key implementations of this method include cascade classifiers that employ LBP (Local Binary Patterns) or Haar-like features. These classifiers are trained using positive examples of a specific object set against random negative images of the same size. Once optimized, the classifier can accurately identify the target object within a specific section of an image. To detect the object throughout the whole image, the classifier systematically examines each segment. It’s essential to differentiate between LBP and Haar-like features. LBP characterizes the local texture of an image by comparing a pixel to its neighboring ones, while Haar-like features measure differences in pixel intensities within neighboring rectangular areas. There are several disadvantages of the sliding window approach, one of them being that it is computationally expensive as a large number of regions need to be classified. To address these issues, region-based methods were invented. The main idea behind these methods was to generate candidate object regions and classify only those regions as containing the object of interest or not.
@@
Another method developed as an improvement over the sliding window approach was the single-stage detection method which has gained popularity in recent years due to its efficiency and good performance. This approach uses a single forward propagation through the network to predict bounding boxes and class probabilities, eliminating the need to generate candidate object regions, and making it faster than region-based approaches. While two-stage detection generates candidate regions in the first stage and refines them in the second stage at the cost of speed and computational efficiency, single-stage detection provides a balance between speed and accuracy by predicting final results in a single pass through the network.
@@
Two-stage detection has been the most widely used approach for detecting wrist abnormalities in recent years. However, there has been limited research on the effectiveness of single-stage detectors in detecting various abnormalities in the wrist, including fractures. In this study, we focus on the effectiveness of SOTA single-stage detectors in detecting wrist abnormalities. Additionally, this study is unique in its use of a large, comprehensively annotated dataset called GRAZPEDWRI-DX presented in a recent publication Nagy et al. (2022). The characteristics and complexity of the dataset are discussed in section 4.
Wrist fractures represent just one of several typical wrist abnormalities, other prevalent conditions include Carpal Tunnel Syndrome (CTS), Ganglion Cysts, Osteoarthritis, Tendinitis, as well as Sprains and Strains. Within the dataset that we use, the distinct objects are categorized as fracture, periostealreaction, metal, pronatorsign, softtissue, bone-anomaly, bonelesion, and foreignbody. It’s crucial to understand that our primary goal is to detect these specific objects rather than diagnose the overarching abnormalities. In our context, the presence of these objects (including fractures) in the wrist can be considered as ’abnormal’. Moreover, the presence of objects other than fractures may suggest another associated wrist abnormality. For instance, soft tissue presence might be indicative of CTS or a ganglion cyst. In CTS, swelling of the synovial tissue that lines the tendons in the carpal tunnel may be observable. Conversely, a ganglion cyst manifests as a soft tissue structure. The term ’bone lesion’ denotes an anomalous area within the bone, severe sprains can involve avulsion fractures where a fragment of bone is pulled away by the ligament.
@@ 1.1 Study Objective & Research Questions
The primary objective of this study is to test the effectiveness of the state-of-the-art YOLO detection models, YOLOv5, YOLOv6, YOLOv7, and YOLOv8 on a comprehensively annotated dataset "GRAZPEDWRI-DX" recently released to the public. We compare the performances of all variants within each YOLO model employed to see whether the use of a compound-scaled version of the same architecture improves its performance. Moreover, this study also investigates how effective these single-stage detection methods are in detecting fractures compared to a two-stage detection method widely used in the past. In addition to conducting object detection across multiple classes, we also evaluate the performance of a conventional CNN in binary classification, specifically in distinguishing between fractures and non-fractures. We hypothesize that fractures in the near vicinity of the wrist in pediatric X-ray images can be detected efficiently using YOLO models proposed by ultralytics (2022), Li et al. (2022), Wang et al. (2022), and ultralytics (2023) respectively.
We analyze the potential of utilizing object detection techniques in answering the following research questions (RQ):
@@
    In the analysis of wrist images, do the single-stage detection models outperform a two-stage detection model widely used in the past?
@@
    Does the use of compound scaled variants within each YOLO algorithm improve its performance in detecting fractures?
  4. 4.
    To what extent can the YOLO surpass conventional CNN architecture and DenseNets in terms of sensitivity in fracture recognition?
@@ 1.2 Contribution
  • A thorough performance assessment of SOTA YOLO detection models on the newly released GRAZPEDWRI-DX dataset, a large and diverse set of pediatric X-ray images. To the best of our knowledge, this is the first study of its kind.
  • An in-depth comparison of the performance of various variants within each YOLO model utilized.
  • Achieved state-of-the-art mean average precision (mAP) score on the GRAZPEDWRI-DX dataset.
  • A detailed performance analysis of single-stage detection models in comparison to the widely-used two-stage detection model, Faster R-CNN.
@@ 3.3 Deep Learning Models For Object Detection
In this study, we employed 4 single-stage detection models, namely YOLOv5, YOLOv6, YOLOv7, and YOLOv8, as well as a two-stage detection model Faster R-CNN. To further optimize the performance of the single-stage models, we experimented with multiple variants of each YOLO model, ranging from 5 to 7 variants. This resulted in a total of 23 wrist abnormality detection procedures.
The YOLO (You Only Look Once) algorithm, initially introduced by Redmon et al. (2015) in 2015, is a single-stage object detection approach that uses a single pass of a convolutional neural network to make predictions about the locations of objects in an image, making it faster than other approaches to date. In 2021, YOLOv4 achieved the highest mean average precision on the MS COCO dataset while also being the fastest real-time object detection algorithm Bochkovskiy et al. (2020). Since its initial release, the algorithm has undergone several improvements, with versions ranging from v1 to v8, with each subsequent version offering smaller volume, higher speed, and higher precision. Fig. 2 illustrates the general structure of YOLO with various backbones used in this study such as CSP, VGG, and EELAN.
@@ 3.3.1 The YOLOv5 Model
The YOLO framework comprises of three components: the backbone, neck, and head. The backbone extracts image features using the CSPDarknet architecture, known for its superior performance Wang et al. (2020). We adopted the same architecture in our research. CSPDarknet involves convolution, pooling, and residual connections represented as:
@@
(Where $F_{i}$ and $F_{i-1}$ are feature maps at $i$-th and $(i-1)$-th layer respectively, $W_{i}$ represents weights and biases, and $f(\cdot)$ applies convolution and pooling operations). The SPP structure is then used to extract multi-scale features from the CSPDarknet’s output:
@@
(Where $F_{SPP}$ denotes multi-scale feature maps, and $g(\cdot)$ performs the SPP operation on $F_{i}$). The neck component adopts the Path Aggregation Network (PANet) to aggregate backbone features, generating higher-level features for output layers. The head constructs output vectors containing class probabilities, objectness scores, and bounding box coordinates. YOLOv5 encompasses five model variants ("n", "s", "m", "l", and "x"), which are compound-scaled versions of the same architecture. These variants offer varying detection accuracy and performance, achieved by adjusting network depth and layer count.
@@ 3.3.2 The YOLOv6 Model
YOLOv6 features an anchor-free design and reparameterized Backbone, with VGG and CSP Backbones used in the "n" and "s" variants, and "m", "l" and "l6" variants respectively. This Backbone is referred to as EfficientRep. The Neck, named Rep-PAN, is similar to YOLOv5, but the Head is efficiently decoupled, improving accuracy and reducing computation by not sharing parameters between the classification and detection branches. The YOLOv6 includes five model variants ("n", "s", "m", "l", and "l6").